Schlumberger is selling for $ 64.91 per share and paid a dividend of $ 1.10 last year. The dividend is expected to grow at 4 percent indefinitely. What is the​ stock's expected rate of​ return?
Answer #1

Ans 5.76%

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
64.91 = 1.1440 / (Ke - 4%)
Ke - 4% = 1.1440 / 64.91
Ke - 4% = 1.76%
Ke = 1.76% + 4%
Ke = 5.76%
D1 = D0* (1 + g)
D1 = 1.10* (1 + 4.00%)
D1 = 1.1440